If the inflation rate is 3% and the nominal interest rate is 8%, what is the...

  1. If the inflation rate is 3% and the nominal interest rate is 8%, what is the real interest rate? Your answer is referred to as the real pre‐tax interest rate.
Answer #1

Based on the Fischer relation,

(1 + Nominal Rate) = (1 + Real Rate) * (1 + Inflation)

(1 + 8%) = (1 + Real rate) * (1 + 3%)

(1 + Real Rate) = 1.08/1.03 = 1.0485

Real Rate = 4.85%